### v2

@@ Abstract
Digital health tools have the potential to significantly improve the delivery of healthcare services. However, their adoption remains comparatively limited due, in part, to challenges surrounding usability and trust. Large Language Models (LLMs) have emerged as general-purpose models with the ability to process complex information and produce human-quality text, presenting a wealth of potential applications in healthcare. Directly applying LLMs in clinical settings is not straightforward, however, with LLMs susceptible to providing inconsistent or nonsensical answers. We describe how LLM-based systems can utilize external tools and provide a novel interface between clinicians and digital technologies. This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLMs in clinical settings such as hallucinations. We illustrate LLM-based interfaces with examples from cardiovascular disease and diabetes risk prediction, highlighting the benefit compared to traditional interfaces for digital tools.
@@
Digital healthcare technologies represent a frontier in medicine and a multitude of tools have been developed sutton2020overview ; dunn2018wearables , in particular risk scores for prognostic modeling moons2009prognosis . Clinical adoption of such methods, however, faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise . One key issue is usability Ratwani2019decade , which can result in errors associated with patient harm Howe2018electronic and contribute to clinician frustration, jeopardizing patient safety Shanafelt2016 ; gardner2019physician .
New tools employing artificial intelligence (AI) and machine learning offer substantial promise Bajwae2021 . Yet these approaches face the same usability challenges as existing digital tools, while AI systems present additional questions around model trust Rajpurkar2022 ; asan2020artificial . These issues have limited the uptake and impact of AI technologies thus far goldfarb2022ai ; Davenport2019 .
@@
Large Language Models (LLMs) offer a potential solution to the challenges faced by digital tools and risk scores if used appropriately. LLMs have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction. Consequently, there is great interest in applying LLMs in healthcare, with potential applications including facilitating clinical documentation, summarizing research papers, or as a chatbot for patients Moor2023 .
Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores. While others have discussed LLMs replacing existing clinical tools, we envisage LLMs forming a novel interface with clinical tools by providing a dynamic output and facilitating deeper interaction with the tools and related information, such as model explainability, medical papers, and clinical guidelines.
@@
In this Comment, we explore a new application of LLMs in healthcare and propose using LLMs to facilitate clinician interactions with AI models and digital tools. Conceptually, this differs substantially from previous applications, such as training medicine-specific LLMs Luo2022 ; Yang2022large or using LLMs for prediction Jiang2023health . Instead, we construct LLM-based systems, offering an intuitive natural language interface that can streamline clinician interactions with multiple digital tools and sources of information, improving efficiency and usability (Fig. 1). By adopting a systems approach, since the LLM does not itself issue predictions or supply specific knowledge, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
@@
LLM-based user interfaces can help engender trust in digital tools and AI models, such as risk scores, by enabling dynamic interactions in the form of natural language dialogues, which has been identified as a key feature for practitioners to deploy machine learning models in healthcare lakkaraju2022rethinking . The most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient. Thus, the ability to adapt and tailor interactions represents a critical advance in the functionality of such tools that can be unlocked by LLMs.
@@
Before Patient
Which features does the risk score use?
Encounter
Why are these features included in the risk score?
How was the risk score validated?
@@
Before Risk
When do clinical guidelines recommend risk scoring?
Scoring
What is the recommended risk score?
@@
During Risk
What is the risk for this patient?
Scoring
What characteristics led to the patient’s risk?
@@
After Risk Scoring
What action is recommended by the guidelines based on the risk?
As a concrete example, let us consider risk scores. Currently, clinicians access these tools via fixed user interfaces or application programming interfaces (APIs). Numerous considerations exist before, during, and after calculating an individual’s prognosis (Table 1), while existing interfaces only serve to calculate risk. Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines, in a manner that specifically addresses their needs or questions without providing superfluous information.
By default, LLMs do not possess the ability to access external tools or information. Methods to extend the functionality of LLMs beyond text generation are in their infancy but already enable powerful LLM-based systems. Below, we describe such a system, illustrated in Fig. 2. Instead of solely relying on an LLM to issue predictions or provide information, we can augment LLMs and enable them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM. This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface. This combines the strengths of LLMs with digital healthcare tools, while overcoming limitations of both, and contrasts previous applications of LLMs in medicine.
1 Illustrative LLM Interfaces
To illustrate the ideas described in this paper, we built example instantiations of LLM-based interfaces for cardiovascular disease (CVD) and diabetes risk prediction (Fig. 3). To demonstrate the versatility of our approach, we took an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbued it with a set of tools using in-context learning, requiring no additional training of the LLM. Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool or accessed specific information and, if so, which tool and with what input. This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source rather than being generated by the LLM. Additional implementation details are provided in Appendix A.
CVD Risk Prediction. We first show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig. 2(a)). We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years. Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD . An example interaction between a physician and the LLM-based system is shown in Fig. 2(a). An illustration of the reasoning process by which the LLM uses external tools is provided in Fig. S.1).
@@ 1 Illustrative LLM Interfaces
In this example, when asked for the patient’s 10-year risk of developing CVD, the LLM uses QRisk3 to estimate the patient’s risk, providing this to the user. The LLM then summarizes the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines. Finally, the LLM uses QRisk3 to recalculate the patient’s counterfactual risk assuming that they were able to reduce their systolic blood pressure to within normal ranges. This allows both the clinician and patient to understand the potential impact of changes to modifiable variables on the patient’s risk. While the clinician could have used the underlying resources to obtain this information, the LLM interface made the interaction simpler and more efficient, which has been identified as a key limitation of digital tools Ratwani2019decade ; Mathews2019 . In the following example, we will explore a more dynamic output extending beyond model use and information retrieval to new functionality.
(a) Cardiovascular disease risk prediction
(b) Diabetes risk prediction
@@
Figure 3: Example interactions with LLM-based interfaces. (a) CVD: The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and provides recommendations for this patient from the NICE clinical guidelines. (b) Diabetes: After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI and understand the underlying methodology.
@@
Diabetes Risk Prediction. We now demonstrate an application for diabetes risk prediction using a machine learning model (Fig. 2(b)). We trained a prognostic model using AutoPrognosis Imrie2023AutoPrognosis , an open-source automated machine learning framework, on a cohort of individuals from UK Biobank Sudlow2015 to estimate a patient’s 10-year risk of developing diabetes. Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, in addition to allowing the LLM access to the model, we provided the LLM with the ability to apply explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions.
@@
As shown in Fig. 2(b), after calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model. In this example, we demonstrate the use of Shapley additive explanations (SHAP) lundberg2017unified . Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions beyond the specific tools and external information sources provided to the LLM. For example, suppose the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified . The LLM is able to explain how this approach works, in a variety of different ways and possibly over multiple interactions with the clinician, allowing specific queries or misunderstandings to be clarified. This demonstrates the utility of LLMs beyond simply using existing tools and illustrates the benefit of their underlying knowledge.

### v3

@@ Abstract
Digital health tools have the potential to significantly improve the delivery of healthcare services. However, their adoption remains comparatively limited due, in part, to challenges surrounding usability and trust. Large Language Models (LLMs) have emerged as general-purpose models with the ability to process complex information and produce human-quality text, presenting a wealth of potential applications in healthcare. Directly applying LLMs in clinical settings is not straightforward, however, with LLMs susceptible to providing inconsistent or nonsensical answers. We demonstrate how LLM-based systems can utilize external tools and provide a novel interface between clinicians and digital technologies. This enhances the utility and practical impact of digital healthcare tools and AI models while addressing current issues with using LLMs in clinical settings such as hallucinations. We illustrate LLM-based interfaces with the example of cardiovascular disease risk prediction. We develop a new prognostic tool using automated machine learning and demonstrate how LLMs can provide a unique interface to both our model and existing risk scores, highlighting the benefit compared to traditional interfaces for digital tools.
1 Introduction
@@
Digital healthcare technologies represent a frontier in medicine. Despite a multitude of tools being developed sutton2020overview ; dunn2018wearables , clinical adoption of such methods faces significant hurdles eichler2007barriers ; Mathews2019 , with some even calling their use “infeasible” Muller2010 and “substantially conceptual” abernethy2022promise . One key issue is usability Ratwani2019decade , which can result in errors associated with patient harm Howe2018electronic and contribute to clinician frustration, jeopardizing patient safety Shanafelt2016 ; gardner2019physician . New tools employing artificial intelligence (AI) and machine learning offer substantial promise, with their impact expected to be felt across all areas of healthcare Bajwae2021 . Yet these approaches face the same usability challenges as existing digital tools, while presenting additional questions around model trust Rajpurkar2022 ; asan2020artificial . Consequently, these issues have limited the uptake and impact of AI technologies in clinical settings thus far goldfarb2022ai ; Davenport2019 ; Kelly2019 .
To improve the usability of clinical predictive models, several approaches have sought to simplify or automate the process of obtaining a prediction. These include points-based scoring systems Gage2001 , web-based calculators Hippisley-Cox2017CVD ; Imrie2023AutoPrognosis , or integration within electronic health records Rothman2013 . While this can make such tools easier to use, simply obtaining a prediction is frequently insufficient and more is required to build model trust with both clinicians Rajpurkar2022 and regulators FDA2019 ; Mourby2021 .
@@
Large Language Models (LLMs) offer a potential solution to the challenges faced by digital tools. LLMs have recently captured the imagination of both the research community and the general public, pushing the boundaries of human-machine interaction. Consequently, there is great interest in applying LLMs in healthcare, with potential applications including facilitating clinical documentation, summarizing research papers, or as a chatbot for patients Moor2023 .
@@
In this paper, we explore a new application of LLMs in healthcare and propose using LLMs to facilitate clinician interactions with AI models and digital tools. Conceptually, this differs substantially from previous applications, such as training medicine-specific LLMs Luo2022 ; Yang2022large or using LLMs for prediction Jiang2023health . Instead, we construct LLM-based systems, offering an intuitive natural language interface that can streamline clinician interactions with multiple digital tools and sources of information, improving efficiency and usability (Fig. 1).
@@
Dynamic interactions in the form of natural language dialogues have been identified as a key feature for practitioners to deploy machine learning models in healthcare lakkaraju2022rethinking . The most appropriate interface or dialogue cannot be pre-specified; instead, it depends on the clinician and the patient. Thus, the ability to adapt and tailor interactions represents a critical advance in the functionality of such tools that can be unlocked by LLMs.
By default, LLMs do not possess the ability to access external tools or information. We augment the base functionality of LLMs and enable them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM. This approach is scalable to multiple predictive models, unifying digital tools within a single, natural language-based interface. By adopting a systems approach, the LLM does not itself issue predictions and can access relevant domain-specific information, rather than needing to supply specific knowledge itself. Consequently, the potential for hallucinations is limited and we ensure actionable information is provided by approved clinical sources.
Figure 1: Clinicians have previously needed to interact directly with digital tools, such as risk scores. While others have discussed LLMs replacing existing clinical tools (1,2), we envisage LLMs forming a novel interface by enabling dynamic interactions and facilitating deeper engagement with tools and related information, such as model explainability, medical papers, and clinical guidelines (3).
As an example of our approach, we specifically focused on risk scoring moons2009prognosis and considered cardiovascular disease (CVD), the most common cause of mortality globally Muthiah2022 . Primary prevention uses prognostic models to estimate the future risk of developing CVD. This allows high-risk individuals to be identified and their risk to be managed via interventions, such as lifestyle modifications or pharmaceuticals. Numerous CVD risk scores have previously been developed, for example the Framingham score DAgostino2008 in the United States, SCORE2 SCORE2 in Europe, and QRisk3 Hippisley-Cox2017CVD in the UK. We begin by developing an improved risk score for CVD using AutoPrognosis 2.0 Imrie2023AutoPrognosis , an open-source automated machine learning framework for developing diagnostic and prognostic models. We then demonstrate a novel LLM-based interface for both our prognostic model and existing risk scores. We provide several illustrative examples of dynamic interactions that substantially extend the capabilities of existing fixed interfaces.
Table 1: Model validation on UK Biobank. C-indices, Brier scores, and Expected/Observed ratios for our approach (AP2), QRisk3, SCORE2, and Framingham score. We report results for the original existing risk scores and the performance following recalibration. Mean performance (95% CI).
C-Index $\uparrow$ Brier Score $\downarrow$ E/O
AP2 0.741 (0.740-0.743) 0.041 (0.041-0.041) 1.003 (1.001-1.006)
Original Risk Scores
Framingham 0.705 (0.703-0.707) 0.047 (0.047-0.047) 2.495 (2.491-2.499)
SCORE2 0.710 (0.709-0.712) 0.042 (0.042-0.042) 1.606 (1.603-1.608)
QRisk3 0.722 (0.721-0.724) 0.044 (0.043-0.044) 1.941 (1.937-1.944)
Recalibrated Risk Scores
Framingham 0.705 (0.703-0.707) 0.041 (0.041-0.041) 1.000 (0.998-1.002)
SCORE2 0.710 (0.709-0.712) 0.041 (0.041-0.041) 1.000 (0.998-1.002)
QRisk3 0.722 (0.721-0.724) 0.041 (0.041-0.041) 1.000 (0.998-1.002)
2 Results
We conducted experiments using data from the UK Biobank Sudlow2015 , a prospective population study of around half a million individuals from the UK, enrolled between 2006 and 2010. UK Biobank collected a broad set of information from participants at enrollment, including health and medical history, blood tests, physical examination, and socio-demographics, with ongoing linkage to healthcare records. The outcome of interest was the incidence of CVD within a 10-year horizon, where CVD was defined as the composite endpoint of myocardial infarction (ICD-10 codes: I21, I22), angina pectoris (I20), stroke (I63, I64), or transient cerebral ischaemic attacks (G45). The cohort contained 407,605 individuals and there were 17,600 CVD events within a 10-year horizon. A flow chart of the study cohort is provided in Fig. S.2 and a summary of patient characteristics is provided in Table S.1. Further details can be found in Section 4.
We used AutoPrognosis 2.0 Imrie2023AutoPrognosis to automatically construct a prognostic model for incidence of CVD. We first validate the performance of our approach compared to existing risk scores and then demonstrate a novel interface for using such risk scores using LLMs.
2.1 Model performance
In Table 1, we compared our prognostic model (AP2) with three existing CVD risk scores, namely QRisk3 Hippisley-Cox2017CVD , SCORE2 SCORE2 , and Framingham score DAgostino2008 . The ensemble model constructed using AutoPrognosis achieved a C-index of 0.741 (95% CI: 0.740-0.743), a Brier score of 0.041 (95% CI: 0.041-0.041), and an expected/observed (E/O) ratio of 1.003 (95% CI: 1.001-1.006), outperforming all three existing risk scores assessed.
(a) Original existing risk scores.
(b) Recalibrated existing risk scores.
Figure 2: Calibration curves. Calibration curves for our approach (AP2), QRisk3, SCORE2, and Framingham score, both before and after calibrating the existing risk scores to the UK Biobank cohort. Observed risk was calculated using Kaplan-Meier estimators kaplan1958nonparametric .
QRisk3 and SCORE2 achieved C-indexes of 0.722 (95% CI: 0.721-0.724) and 0.710 (95% CI: 0.709-0.712), respectively, with similar Brier scores (QRisk3: 0.044, 95% CI: 0.043-0.044; SCORE2: 0.042, 95% CI: 0.042-0.42). Both approaches overpredicted risk in the UK Biobank population (E/O ratio QRisk3: 1.941, 95% CI: 1.937-1.944; SCORE2: 1.606, 95% CI 1.603-1.608) consistent with previous findings Parsons2023 . The Framingham score was the weakest performing risk score, achieving a C-index of 0.705 (95% CI: 0.703-0.707). This underperformance likely was a consequence of the Framingham score being developed on a substantially different population from the USA.
Calibration curves (Fig. 2) show that AutoPrognosis predictions were well calibrated across all risk deciles. In contrast, the existing risk scores overpredicted risk for all individuals in the UK Biobank cohort (Fig. 2(a)). We recalibrated the existing risk scores by scaling the predictions according to the overall incidence in the UK Biobank cohort and the mean predicted risk (Fig. 2(b)). Following recalibration, QRisk3 exhibited good calibration across all risk deciles, while SCORE2 and Framingham score both modestly overpredicted risk in the lower risk deciles and underpredicted risk in the highest risk decile.
(a) Original existing risk scores.
(b) Recalibrated existing risk scores.
Figure 3: Decision curve analysis. Our approach (AP2) provides greater net benefit at all thresholds than the existing risk scores (Framingham score, SCORE2, and QRisk3) and the baseline strategies (All and None).
Figure 4: Feature importance. SHAP values of the variables included in the AutoPrognosis model.
To assess model performance in a manner that considers the effect on clinical decision making, we additionally performed decision curve analysis Vickers2006 ; Vickers2008 , as recommended in the TRIPOD guidelines Moons2015 . Decision curve analysis is a method to evaluate the clinical utility of a model by assessing the net benefit for a range of risk thresholds at which a decision maker would act or intervene. The model with the higher net benefit at any given threshold is preferred. We compared our model with the existing risk scores, as well as baseline strategies that assume all patients will develop CVD (All) or none will (None). Our model achieved a greater net benefit at all decision thresholds compared to the existing risk scores (Fig. 3). Net benefit takes into account both discrimination and calibration. While recalibrating the existing risk scores improved their net benefit, the AutoPrognosis model still resulted in the highest net benefit at all thresholds.
To better understand the rationale for the predictions and drivers of the outperformance of our approach, we assessed feature importance by calculating SHAP values lundberg2017unified . SHAP values for the 20 features included in the final version of our model are shown in Fig. 4. Consistent with clinical knowledge, age, sex, and blood pressure are three of the most important features and were included in all three existing risk scores considered in this work. However, a number of additional features, including both laboratory and nonlaboratory tests, also significantly contributed to model predictions. A number of these features have been shown to be risk factors for CVD, but have not been incorporated into other risk scores. For example, Apolipoprotein B (ApoB) is a primary component of several lipoproteins, including low-density lipoprotein (LDL), and transports cholesterol and other lipids in the blood. High levels of ApoB are the primary driver of atherosclerosis and there is evidence that ApoB is a more accurate predictor of CVD than total cholesterol or LDL Marston2022 ; Behbodikhah2021 . ApoB is deemed one of the most important variables in our approach (Fig. 4), along with Apolipoprotein A and Lipoprotein A which have also been shown to be risk factors for CVD Erqou2010 . Finally, “Overall health rating”, a non-laboratory, self-reported feature, was found to be predictive, mirroring the findings of Alaa et al. regarding the benefits of non-laboratory variables for CVD risk prediction Alaa2019 .
@@ 2.1 Model performance
Before Which features does the risk score use?
Patient Why are these features included in the risk score?
Encounter How was the risk score validated?
@@
Before Risk When do clinical guidelines recommend risk scoring?
Scoring What is the recommended risk score?
@@
During Risk What is the risk for this patient?
Scoring What characteristics led to the patient’s risk?
@@
After Risk Scoring What action is recommended by the guidelines based on the risk?
@@
2.2 LLM-based Interfaces
While LLMs are powerful models for natural language processing, by default, they do not possess the ability to access external tools or information. Methods to extend the functionality of LLMs beyond text generation are in their infancy but can already be used to significantly expand the capabilities of LLMs schick2023toolformer ; nakano2021webgpt .
In this section, we construct LLM-based systems and demonstrate multiple examples of how such systems can provide a novel interface for digital health tools, in particular clinical risk scores. The LLM-based systems can incorporate numerous external tools, sources of information, and clinical data (illustrated in Fig. 5). Instead of using an LLM to issue predictions directly, we augmented LLMs and enabled them to access approved medical tools and other sources of information, thereby not solely relying on the inherent capabilities of a given LLM, while unifying multiple digital tools within a single, natural language-based interface.
To demonstrate the versatility of our approach, we took an off-the-shelf pretrained LLM, GPT-4 openai2023gpt4 , and imbued it with a set of tools using in-context learning, requiring no additional training of the LLM. Additionally, we implemented a “source” functionality that allows the user to see whether the LLM used a tool or accessed specific information and, if so, which tool and with what input. This helps avoid hallucinations since it enables verification that the information was issued by an underlying clinical tool or source rather than being generated by the LLM. Additional implementation details are provided in Section 4.
@@ 2.2 LLM-based Interfaces
Figure 6: Example interactions with LLM-based interfaces. (a) Clinicians can ask the LLM questions about the AutoPrognosis risk score, such as what variables are included, the underlying methodology, and the cohort on which the model was derived. (b) After calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using explainable AI to improve understanding of the model predictions.
AutoPrognosis. We constructed an LLM-based interface for the machine learning model developed in this work (Fig. 6). In our first example, we demonstrate the role such an interface could have upon model deployment (Fig. 6(a). A clinician might have many questions about a risk score before using the predictions to inform recommendations or interventions (Table 2). Here, the LLM provided specific information about the variables used by the risk score, the underlying methodology, how the model was constructed, and the cohort of individuals used to train the model.
@@
Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, in addition to allowing the LLM to access the model, we provided the LLM with the ability to apply methods from explainable AI (XAI) Imrie2023 to the underlying model, allowing users to investigate the rationale for predictions.
Figure 7: Example interaction with an LLM-based interface for QRisk3. The LLM uses QRisk3 to calculate the patient’s 10-year risk of developing CVD, explains why certain features are included in the risk score using the QRisk3 paper Hippisley-Cox2017CVD , and provides recommendations for this patient from the NICE clinical guidelines.
@@
As shown in Fig. 6(b), after calculating the patient’s risk using the AutoPrognosis model, the clinician can query why this prediction was issued using XAI, facilitating a deeper understanding of the model. In this example, we used Shapley additive explanations (SHAP) lundberg2017unified to explain the prediction. For this individual, their estimated 4.2% risk was primarily caused by their age and family history of CVD, mitigated by being a woman, not smoking, and low levels of lipoprotein (a). An additional interaction can be found in Fig. S.3. Due to their general pretraining, LLMs have knowledge of many topics; this can provide valuable additional information during interactions beyond the specific tools and external information sources provided to the LLM. For example, suppose the clinician is not familiar with the underlying XAI methodology, SHAP lundberg2017unified . The LLM could explain how this approach works, in a variety of different ways and possibly over multiple interactions with the clinician, allowing specific queries or misunderstandings to be clarified. This demonstrates the utility of LLMs beyond simply using existing tools and the benefit of their underlying knowledge.
QRisk3. We also show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig. 7). We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years. Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD . An example interaction between a physician and the LLM-based system is shown in Fig. 7. An illustration of the reasoning process by which the LLM uses external tools is provided in Fig. S.1.
In this example, when asked for the patient’s 10-year risk of developing CVD, the LLM used QRisk3 to estimate the patient’s risk, providing this to the user. The LLM then summarized the QRisk3 paper to explain the inclusion of certain features before providing the recommended action for this patient from the NICE clinical guidelines. Finally, the LLM used QRisk3 to recalculate the patient’s counterfactual risk assuming that they were able to reduce their systolic blood pressure to within normal ranges. This allows both the clinician and patient to understand the potential impact of changes to modifiable variables on the patient’s risk. While the clinician could have used the underlying resources to obtain this information, the LLM interface made the interaction simpler and more efficient, which has been identified as a key limitation of digital tools Ratwani2019decade ; Mathews2019 .